OpenStudy (anonymous):

A painter leans a 20-ft ladder against a building. The base of the ladder is 12 ft from the building. To the nearest foot, how high on the building does the ladder reach?

OpenStudy (johnweldon1993):

We would need an angle to work with here...

OpenStudy (anonymous):

it was no angle @johnweldon1993

OpenStudy (johnweldon1993):

Hmm.... Lets find the angle then....we have a side adjacent to the angle and the hypotenuse...what is the arccos of 12/20 That is 53.1301024 degrees...(lets say 53 degrees) Now we have an angle....a hypotenuse...and we want the side opposite that angle...so sin(53.1301024) = o / 20 Multiply both sides by 20 o = 20 * sin(53.1301024) o = 16 feet...now lets check using the good old pythagorean theorem a^2 + b^2 = c^2 12^2 + 16^2 = 20^2 144 + 256 = 400 400 = 400 That is correct...so it is 16 feet high

OpenStudy (anonymous):

thanx @johnweldon1993

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

In , LM = 11 and . Find KL. Leave your answer in simplest radical form.

OpenStudy (anonymous):

@johnweldon1993 can you help me

OpenStudy (johnweldon1993):

I need 1 more side...all you write was that LM = 11 ...I need 1 more side...

OpenStudy (anonymous):

m<k=45

OpenStudy (johnweldon1993):

Oh okay...well then we know that a 45 45 90 triangle has the ratio 1 : 1 : √2 So we know that 1 leg is 11.....so the other leg must be 11....and the hypotenuse KL would be 11 times √2 11√2 would be your answer

OpenStudy (anonymous):

thank you